### 16 Dec 2023 v3

# Enhancing Accuracy and Robustness of Steering Angle Prediction with Attention Mechanism

Abstract
In this paper, we investigate the two most popular families of deep neural architectures (i.e., ResNets and InceptionNets) for the autonomous driving task of steering angle prediction. To ensure a comprehensive comparison, we conducted experiments on the Kaggle SAP dataset and custom dataset and carefully examined a range of different model sizes within both the ResNet and InceptionNet families. Our derived models can achieve state-of-the-art results in terms of steering angle MSE. In addition to this analysis, we introduced the attention mechanism to enhance steering angle prediction. This attention mechanism facilitated an in-depth exploration of the model’s selective focus on essential elements within the input data. Furthermore, recognizing the importance of security and robustness in autonomous driving assessed the resilience of our models to adversarial attacks.
Index Terms:
Steering-angle, InceptionNet, ResNet, Attention,adversarial attacks
I Introduction
In the era of autonomous vehicles, the precise prediction of steering angles is essential in ensuring safe and efficient self-driving operations of the vehicle. The accuracy of this steering angle prediction directly impacts the vehicle’s ability to navigate complex and dynamic environments. Deep learning models have emerged as powerful tools for tackling this challenge, striving to provide real-time and reliable steering angle predictions [1]. This paper embarks on a comprehensive exploration of compact InceptionNet and ResNet model variations, with a specific focus on their effectiveness in this crucial task. Additionally, we investigate the incorporation of attention mechanisms, which improve the accuracy of these models by allowing selective focus on essential input data elements [2]. Next, We will show that our attention mechanism plays an important role in improving the clarity and precision of predictions for steering angles, enabling vehicles to make better-informed choices while driving.
Although deep learning models show significant potential across different applications, it’s important to acknowledge their susceptibility to adversarial attacks. Prior studies have highlighted the presence of adversarial examples in the realm of image classification. These examples involve making tiny, often unnoticeable changes to input data, which can lead to substantial changes in the model’s output. This inherent susceptibility raises concerns about the safety and reliability of neural networks, particularly in critical fields like self-driving cars. In the context of autonomous vehicles, where decision-making heavily relies on neural networks, the potential impact of adversarial attacks is of paramount importance. The ability of an attacker to subtly manipulate input data and affect a self-driving car’s behavior poses a substantial risk to road safety and system dependability. Our study also explores how our attention mechanism might counteract these models’ susceptibility to adversarial attacks. We aim to investigate whether incorporating attention mechanisms enhances the models’ resilience to security threats.Additionally, we explore potential variations in how the models handle these tasks. This thorough analysis will provide insights into how secure these models are and contribute to the safety and reliability of autonomous driving systems.
II Related Work
Autonomous driving is a rapidly evolving field with a primary focus on enhancing the safety, efficiency, and accuracy of self-driving vehicles. Researchers have been dedicated to developing neural network architectures that are not only highly accurate but also computationally efficient, catering to the demands of real-time autonomous navigation. One specific and critical task in autonomous driving is steering angle prediction, a fundamental component of a vehicle’s control system. Steering angle prediction involves training artificial intelligence models to anticipate the appropriate steering angle based on road curvature and the vehicle’s environmental surroundings. This predictive capability is important for autonomous vehicles to navigate safely [3].
The current state of the work done in steering angle prediction is early compared to other areas like image recognition. Steering angle prediction continues to rely more heavily on visual processing and understanding. Currently, even Tesla has begun to use pure vision in their cars [4]. [5] survey different deep learning (DL) methods for steering angle prediction. The goal of this research is to explore the steering angle control of autonomous vehicles (AV) using the two most popular deep learning architectures (i.e., InceptionNet and ResNet models). Needless to say, researchers tend to focus on steering angle prediction using deep convolutional neural networks, which is not surprising given their success in other computer vision tasks.
Steering angle prediction needs to be accurate, fast, and efficient. If the prediction is slow, it could cause life-threatening harm to a pedestrian. Since steering angle prediction is an active topic, there is not a considerable amount of recent work done in this area. While there are continuous changes in the field, the ResNet architecture remains a key player in steering angle prediction [6; 7; 8]. For example, [6] surveyed different methods for achieving steering angle prediction. They begin by categorizing previous works on steering angle calculations into two major approaches. One is traditional computer vision-based and the other is neural network-based. Most of the approaches rely on cameras, sensors, and/or radar. The results of this research compilation showed that the combination of ResNet50 deep architecture with event cameras gives a better prediction of the due wheel angle.
However the cost to train and deploy ResNet50 is high. This is mainly due to the large number of parameters in a ResNet50 model. Other research has fairly large and complex models for pushing the boundaries of autonomous driving. [1] investigate creating a new end-to-end architecture design for speed and steering angle prediction. This architecture design consists of creating layers within a stack of blocks and connecting them like a puzzle [1]. It is compared to other end-to-end architectures used in steering angle prediction. The results of this research show that these new end-to-end models have better optimization from the use of additional data points, like speed, causing these models to perform well with a reduced number of parameters.
The ResNet architecture is not the only architecture for steering angle predictions. ResNet research does appear to be more popular in the literature, even though InceptionNet is another architecture that can address steering angle predictions. [9] found InceptionNet as the best model for various autonomous driving tasks. [10] found that InceptionNet extracted qualitatively equivalent information as its ResNet counterpart. This paper explores both ResNets and InceptionNets at the compact end of the complexity spectrum.
The attention mechanism has been used in a variety of computer vision tasks. [2] proposed a residual attention network to focus on an image’s most informative parts, allowing for more efficient and accurate classification. The attention mechanism works by adaptively weighing different features of the image, based on the context and spatial location. The attention mechanism includes the development of frameworks that incorporate both bottom-up and top-down structures in order to extract global features. At lower resolutions, the bottom-up feed-forward structure generates feature maps with robust semantic information, whereas the top-down network generates dense features for inference on each pixel. The present state-of-the-art performance in image segmentation has been attained by inserting skip connections between the bottom and top feature maps [11]. The study demonstrated the effectiveness of the attention mechanism in improving the network’s performance. Using an attention-based neural network architecture, [12]proposed a method to reconstruct dense point clouds from a single image that can accurately capture the geometry and texture of the input image. The attention module is configured to learn a weighting method that weights each point in the point cloud based on its importance to the input image. During training, the weighting scheme is learned by minimizing a reconstruction loss that measures the difference between the rebuilt point cloud to the ground truth point cloud.
With the progress in autonomous vehicle technology, the robustness of the underlying neural network models becomes paramount. Ensuring the reliability and safety of self-driving cars demands an understanding of how these models perform under adversarial conditions, such as deliberate input perturbations aimed at causing mispredictions. While adversarial attacks on neural networks have been extensively explored in different fields, mainly in image classification, the use of these attacks in autonomous driving systems, especially regarding steering angle prediction, is becoming an increasingly significant area of focus [13]. In this paper, we evaluate the vulnerability of the neural networks particularly in the context of steering angle prediction, to adversarial attacks such as Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD). Our goal is to improve the overall safety and security of autonomous driving systems.
Figure 1: Illustration of deep-learning-based steering angle prediction on visual data
III Methodology
A general description of supervised training is having a CNN model take different images and labels to learn what to look for. The learned weights are how the algorithm is able to differentiate different aspects of an image. In our case, the algorithm needs to learn different turns and report the corresponding turning angles of the steering wheel (in degrees). These roads must vary to best test the models and evaluate outcomes for accuracy. To address the inefficiency issue faced with training and running large CNN models for steering angle prediction, more compact models must be considered. The first goal is to find a compact model with an accuracy score that does as well as or better than its larger counterparts. This paper explores variations of compact ResNet and InceptionNet models.
We present the outcomes of various sizes for each of the two architects. For example, we built and trained eight ResNet architectures; ResNet20, ResNet22, ResNet24, ResNet26, ResNet28, ResNet30, ResNet32, and ResNet34. We built seven InceptionNet models to match similar numbers of parameters so that we could compare the architectures at different complexity levels. The number of parameters and block layers for each of the ResNet and InceptionNet models is illustrated in Table I and II. We utilized pre-trained ResNet and Inception models from the PyTorch library as baseline models before building our own ResNet and InceptionNet models. Figure 1 offers a high-level illustration of using deep learning models and visual data for steering angle prediction.
For evaluating the different architectures, mean squared error (MSE) loss was used. We will begin by providing an overview of the deep learning models employed in our experiments, which include compact versions of InceptionNet and ResNet architectures. Subsequently, we will introduce an attention mechanism designed to enhance the precision and interpretability of steering angle predictions. Furthermore, we’ll examine how well these models stand up to adversarial attacks, including the Fast Gradient Sign Method(FGSM) and Projected Gradient Descent(PGD) to enhance the safety and dependability of self-driving car systems.
III-A ResNet
Preliminary experiments included running five pre-trained ResNet models. The largest two models (101 & 152) were too large and unusable for our purposes. PyTorch’s pre-trained ResNet18 and ResNet34 performed very poorly when compared to the pre-trained ResNet50, plus we needed varying degrees between the two models. This was the starting point but we needed to produce our own ResNet models to help tease out any nuances between the architectures. We constructed our ResNet models based on the fundamental ResNet 18 and ResNet 34 basic block methodology based on the original ResNet paper. The input layer of ResNet models takes in an image of dimension 224 by 224. For those interested in ResNet architecture details, please see [14].
TABLE I: ResNet Model Parameters & Block Layers
ResNet Model Parameters Block Layers
ResNet20 12.8 million (2,2,3,2)
ResNet22 13.1 million (2,3,3,2)
ResNet24 14.3 million (2,3,4,2)
ResNet26 17.9 million (3,3,3,3)
ResNet28 19.1 million (3,3,4,3)
ResNet30 19.4 million (3,4,4,3)
ResNet32 20.6 million (3,4,5,3)
ResNet34 21.7 million (3,4,6,3)
TABLE II: InceptionNet Model Parameters & Block Layers
InceptionNet Model Parameters Block Layers
InceptionNet 13.0 million (2,5,2)
InceptionNet a 14.5 million (5,5,2)
InceptionNet b 15.5 million (2,8,2)
InceptionNet c 17.0 million (5,8,2)
InceptionNet d 17.6 million (2,5,5)
InceptionNet e 20.2 million (2,8,5)
InceptionNet f 21.7 million (5,8,5)
III-B InceptionNet
Traditional InceptionNet is different from traditional ResNet. The framework of an InceptionNet model aims to be powerful and efficient as it combines the strength of various filter sizes (e.g., 1x1, 3x3, 5x5, and max pooling) [15]. Another important difference is that InceptionNet concatenates while ResNet summates at the end of blocks. Similar to ResNets, InceptionNets can be adjusted in terms of the number of layers or parameters to assess and enhance the accuracy. The Inception block uses various filters to convolve the images and then concatenates the results.
Within a certain range, we attempted to add accuracy by increasing the size of the blocks in each of the main architectural sections, which in turn changes the number of filters throughout the layout. These changes added to the level of complexity within the InceptionNet blocks. For further interest in the InceptionNet Block architecture, please see [15].
III-C Attention-aware Deep Model for Steering Angle Prediction
Figure 2: Comparison of ResNet32 and ResNet32 with attention modules (Same size of residual unit)
In this section, inspired by [2], we propose to add the attention mechanism for the task of steering angle prediction. As an example, we take ResNet32 as the baseline model for our experiments. Figure 2 shows a comparison of our basic ResNet32 network and ResNet32 network with the addition of attention modules, where both networks have the same number of residual units to ensure a fair comparison. As shown in figure 2 the attention module consists of two basic branches. The trunk branch and the attention mask branch. The trunk branch works as a standard encoder to extract features from the input. On the other hand, the attention mask branch learns a mask of the same size as the extracted features from the encoder on which the implementation of an effective attention mechanism relies. In order to extract the global features, the input features undergo several down-samplings. These global features are then extended to their original sizes through symmetrical up-sampling to obtain the mask. By calculating the dot product of the mask and the output of the trunk branch, the mask then escorts output features at each point.
The attention module includes three attention blocks, each of which consists of three operations: the feature fusion operation, the feature selection operation, and the feature amplification operation. The feature selection operation separates the useful features from the unimportant ones while the feature fusion operation combines the feature maps from the trunk network and the preceding attention block. The operation of feature amplification amplifies the crucial features to improve their representation.
III-D Model Robustness Assessment
We evaluate how adversarial attacks impact the ResNet32 model, both with and without the attention mechanism. To ensure fair comparisons, each network is configured with the same number of residual units. The attention module, as explained previously, incorporates trunk and attention mask branches, facilitating an effective attention mechanism that selectively emphasizes essential features in the input data.
In addition to enhancing accuracy through attention mechanisms,our research also delves into the model’s ability to resist adversarial attacks, such as Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD). These attacks introduce subtle perturbations to input data and are applied to both version of the ResNet32 model- with and without attention. Through these adversarial tests, we aim to assess whether the attention mechanism enhances the model’s security and its capacity to cope with adversarial manipulations [16].
IV Experimental Setup
IV-A Testing Environment
Experiments were run on the Ohio Supercomputer Center’s Pitzer cluster which includes dual NVIDIA Tesla V100 GPUs with 32GB memory [17]. The deep learning framework used was PyTorch. All imports will be based on this framework using Python 3.
IV-B Evaluation
We focused on the validation loss and logged those scores using the Tensorboard. We use mean squared error (MSE) as the loss function:
$MSE=\frac{1}{n}\displaystyle\sum_{n=1}^{n}(r_{i}-\hat{r_{i}})^{2}\,,$ (1)
where $r_{i}$ is actual steering angle (in radians) and $\hat{r_{i}}$ is predicted steering angle (in radians). We implemented an early stopping mechanism, terminating model training if no improvement was observed after 10 epochs.
Each model was trained one at a time and its checkpoint was saved such that it could be included in a visualization using Tensorboard. Models were run for 50 epochs and the batch size tested for all models varied from 16 to 128 in search of the best outcome. Generally, batch sizes of 32 were found to have the best results. Epochs of at least 50 were needed as the ResNet model’s validation loss score decreased earlier in their epochs, however, the InceptionNet models performed more similarly after 30 or more epochs. This need for at least 30 or so epochs was a recurring theme when comparing Inception and Residual architects at all levels of complexity. Figure 3 highlights the quick drop in validation loss scores in earlier epochs, which was a common observation in models for the ResNet architecture.
Figure 3: Magenta line represents ResNet20, & Orange line represents InceptionNet.
V Results
V-A Quantitative
Our first analysis was with a publicly available Kaggle data set of approximately 97,330 images [18]. In this paper, we refer to this dataset as the Kaggle Steering Angle Prediction dataset or simply the Kaggle SAP dataset. If we focus solely on MSE, the ResNet32 model achieves the lowest error at 0.05293. (see Table III). This is a competitive score as our ResNet model performed similarly (or better) than that of recent work [19; 20] [^1] We do even better in our second analysis.. The InceptionNet architecture had a score of 0.05846 as its lowest. However, this does not tell the whole story. Figure 4 plots all of the tested models with their complexity measure on the x-axis and their MSE scores on the y-axis. The plotted lines intersect around 17.9 or 18 million parameters.
TABLE III: MSE Scores on the Kaggle SAP dataset
ResNet Model MSE InceptionNet Model MSE
ResNet20 0.06557 InceptionNet 0.06044
ResNet22 0.06371 InceptionNet a 0.05945
ResNet24 0.06721 InceptionNet b 0.05849
ResNet26 0.05832 InceptionNet c 0.05846
ResNet28 0.05552 InceptionNet d 0.05961
ResNet30 0.05641 InceptionNet e 0.05802
ResNet32 0.05293 InceptionNet f 0.05907
ResNet34 0.05425
12121212131313131414141415151515161616161717171718181818191919192020202021212121222222225.255.255.255.255.55.55.55.55.755.755.755.7566666.256.256.256.256.56.56.56.56.756.756.756.75No. of Parameters (in millions)MSE Loss $(10^{-2})$InceptionNetResNet
Figure 4: MSE Loss vs Number of Parameters on the Kaggle SAP dataset
This is approximately where the two architectures are equivalent. Thus, with our main idea to have smaller neural networks, we can see that at a smaller complexity level, the InceptionNet clearly outperforms the ResNet models. We see that at deeper and more complex models the ResNet architecture outperforms the InceptionNet models after about 18 million parameters.
To make sure that the results were not just an anomaly of the Kaggle SAP dataset, we ran all of the models on a second dataset. Our second analysis included a dataset of approximately 10,765 images. We refer to this dataset as the custom dataset. The images can be found here [21]. This dataset was created by the authors using the Udacity self-driving-car simulator [22]. Table IV highlights the best performing model is our InceptionNet b model. Interestingly, all but one of our InceptionNet models outperformed its parameter-comparable ResNet counterpart. So not only do we see the best performance from the InceptionNet architecture, but we also see the InceptionNet architecture outperforming ResNet at various ranges. More importantly, we see InceptionNet perform very well at the lower-end and mid-range-sized models. Figure 5 compares the MSE loss of ResNet and Inception Net models, showing optimal performance with InceptionNet around 17 million parameters.
While the InceptionNet architecture demonstrated good results on the custom dataset, our choice to utilize ResNet32 for the Kaggle SAP dataset was informed by the dataset’s unique characteristics. The Kaggle dataset, being substantially larger in scale, required models capable of handling a larger volume of data effectively. In this context, ResNet32 and ResNet26 consistently outperformed InceptionNet, demonstrating both accuracy and efficiency. Therefore, we chose models based on the specific needs of each dataset to ensure strong performance and accurate steering angle predictions. The subsequent integration of attention modules was strategically employed to further enhance the chosen ResNet models, ultimately reducing MSE and delivering superior results across both datasets. Our objective was to explore how adding attention mechanism might boost model efficiency in lowering MSE, and our findings highlight the effectiveness of this approach for enhancing steering angle prediction accuracy. To this end, we conducted a comparison between the baseline ResNet32 model and ResNet32 with attention modules, utilizing the Kaggle SAP dataset. The results depicted in Figure 6, reveal a significant reduction in MSE during validation after incorporating attention modules into the ResNet32 model at around 70 epochs during training. The loss of our model with attention modules smoothly goes down throughout the whole training process while we notice some sudden bumps around 60 to 80 epochs for ResNet32 without attention modules. Additionally, we evaluated the performance of the ResNet26 model with and without attention modules, as ResNet26 exhibited the best results for the custom images. Figure 7 shows a comparison between the mean squared error (MSE) scores of two models during validation: the baseline ResNet26 model and the ResNet26 model with integrated attention modules. The MSE scores are plotted against the number of epochs. According to Figure 7, the MSE of the attention-aware ResNet26 model is significantly decreased with the number of epochs. Except for epoch 70, where the performance of both models was similar, attention modules improved the performance of the ResNet26 baseline model. After 70 epochs, our model began to decay smoothly, while we could still see some bumps in the MSE score for ResNet26 without attention modules. At epoch 100, we observe a 6.83% reduction in MSE for the ResNet32 on the Kaggle SAP dataset and a 6.09% reduction for the ResNet26 on the custom dataset when the attention mechanism was employed. Figure 8 illustrates the percentage reduction in MSE at 100 epochs during validation. These findings show that the integration of attention modules into the ResNets can yield a lower MSE.
TABLE IV: MSE Scores on Custom Dataset
ResNet Model MSE InceptionNet Model MSE
ResNet20 0.05527 InceptionNet 0.04886
ResNet22 0.05081 InceptionNet a 0.05692
ResNet24 0.04701 InceptionNet b 0.03828
ResNet26 0.04642 InceptionNet c 0.04395
ResNet28 0.04774 InceptionNet d 0.03981
ResNet30 0.05311 InceptionNet e 0.04402
ResNet32 0.04817 InceptionNet f 0.04557
ResNet34 0.04646
12121212131313131414141415151515161616161717171718181818191919192020202021212121222222223.753.753.753.7544444.254.254.254.254.54.54.54.54.754.754.754.7555555.255.255.255.255.55.55.55.55.755.755.755.75No. of Parameters (in millions)MSE Loss $(10^{-2})$InceptionNetResNet
Figure 5: MSE Loss vs Number of Parameters for the Custom dataset
Continuing from the selection of the ResNet32 model for our attention experiments, we also leveraged the same model for our subsequent adversarial robustness investigations. By using the ResNet32 model, we aimed to understand how integrating attention modules influenced its response to adversarial attacks. These attacks aimed to test the model’s susceptibility and its capability to handle alterations in the input data . We employed two well-established adversarial attack techniques: the Fast Gradient Sign Method (FGSM)[16] and Projected Gradient Descent (PGD) [23]. PGD differs from FGSM as it is an iterative approach, involving multiple steps to iteratively refine adversarial examples. The choice of ’Epsilon (Eps)’ values in our experiments determined the magnitude of these adversarial perturbations, influencing the severity of the attacks [24]. The results of our adversarial attack experiments, including the MSE scores for the ResNet32 model with and without integrated attention modules, are presented in Table V. These results offer crucial understanding of the model’s security and its resilience to adversarial threats, especially in autonomous vehicle systems. It’s essential to understand that the application of adversarial attacks in autonomous vehicles is a pressing concern, given the potential implications for road safety and system reliability. While previous research has indicated the vulnerability of neural networks to such attacks, our study takes a step further by being the first to test the adversarial robustness of attention-equipped ResNets for steering angle prediction.
While it’s well-documented that adversarial attacks can lead to an increase in MSE loss, our research delves into whether the integration of attention modules can act as a robust countermeasure. In our experiments, we observe how the model’s performance evolves when subjected to adversarial attacks. While the MSE loss does indeed increase, our results indicate that, with the inclusion of attention modules, the degree of this increase is notably mitigated as shown in Table V.
TABLE V: Adversarial attack results
Eps FGSM with attn FGSMw/o attn PGD with attn PGDw/o attn
0.01 0.200 0.214 2.581 4.763
3 0.291 0.336 2.695 5.853
V-B Qualitative
To better understand the differences between the different types of architectures in our experiments, we looked into the saliency maps of the models. Saliency maps provide insight into how the three architectures differ in their focus on an image (Fig. 9). The image used for the saliency example was pre-processed. We calculated the model gradients and subsequently determined the maximum absolute values across the image channels for each pixel. This process was employed to emphasize the contribution of the output in generating the saliency maps. The saliency images show primarily three colors; black which indicates no focus or attention; red which represents some focus or attention; and yellow which denotes the highest level of focus or attention. Our analysis is still very preliminary; nonetheless, it appears ResNet and ResNet with attention modules are more general in their viewing of an image, while InceptionNet appears to be more focused on the road end within the image.
1010101020202020303030304040404050505050606060607070707080808080909090901001001001004.54.54.54.555555.55.55.55.566666.56.56.56.577777.57.57.57.58888No. of epochsMSE Loss $(10^{-2})$ResNet32 w attentionResNet32
Figure 6: MSE comparison between ResNet32 vs ResNet32 with attention modules for the Kaggle SAP dataset
1010101020202020303030304040404050505050606060607070707080808080909090901001001001004.24.24.24.24.64.64.64.655555.45.45.45.45.85.85.85.8No. of epochsMSE Loss $(10^{-2})$ResNet26 w attentionResNet26
Figure 7: MSE comparison between ResNet26 vs ResNet26 with attention modules for the Custom dataset
Figure 8: Performance improvement of ResNet-26 and ResNet-32 with attention modules added
(a) Original image
(b) ResNet32
(c) Inception d
(d) ResNet32 with attention
Figure 9: Saliency maps representing the focus areas of the different models. The saliency maps use three colors. Black signifies no focus, red indicates some focus, and yellow indicates the highest level of focus. To generate a saliency map we get a predicted output from the model. We take the max score of this output and perform back-propagation on the max score with respect to the original input. We use layer 4 saliency as an example. For the last three subfigures, the saliency maps are blended with the original image with a blending ratio of 0.75.
VI Conclusion and Future Work
This research first examines the two main veins of modern neural network architectures and their intra-variations. The models explored for this research include various InceptionNet and ResNet depths of complexity. If there is no limit on the model/data size, then it appears that the ResNet architecture tends to perform better with deeper and more complex structures. In addition, we have also added attention modules to ResNet architectures, which has further increased the models’ performance by decreasing the MSE up to 6.83% for the SAP dataset and up to 6.09% for the custom dataset. All of the programs and the custom dataset used in this study are publicly accessible in this repository [21]. Our investigation extended into the realm of adversarial robustness, where we systematically tested the resilience of our models against adversarial attacks, thereby uncovering potential vulnerabilities. Intriguingly, the incorporation of attention modules demonstrated a promising capability to fortify the models against adversarial attacks, thereby reducing their adverse impact on predictive accuracy. In our future work, we will explore different defense mechanisms to overcome these adversarial attacks.